OpenStudy (anonymous):

Yara rolls a standard die and flips a coin. What is the probability that she gets an even number and heads?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

sample space has 6*2=12 elements. you want these (2,h) (4,h) (6,h) so \[\frac{3}{12}=\frac{1}{4}\]

OpenStudy (anonymous):

both 1/2

OpenStudy (anonymous):

oh heck no. even number AND heads. i wrote them all out

OpenStudy (anonymous):

why yar 3/6=1/2 isnt it?

OpenStudy (anonymous):

(0.5)(0.5)=0.25 , as satellite sed